# Inside the black box: Neural network-based real-time prediction of US recessions

### v2

@@ Abstract
A standard feedforward neural network (FFN) and two specific types of recurrent neural networks, long short-term memory (LSTM) and gated recurrent unit (GRU), are used for modeling US recessions in the period from 1967 to 2021. The estimated models are then employed to conduct real-time predictions of the Great Recession and the Covid-19 recession in the US. Their predictive performances are compared to those of the traditional linear models, the standard logit model and the ridge logit model. The out-of-sample performance suggests the application of LSTM and GRU in the area of recession forecasting, especially for the long-term forecasting tasks. They outperform other types of models across five different forecast horizons with respect to a selected set of statistical metrics. Shapley additive explanations (SHAP) method is applied to GRU and the ridge logit model as the best performer in the neural network and linear model group, respectively, to gain insight into the variable importance. The evaluation of variable importance differs between GRU and the ridge logit model, as reflected in their unequal variable orders determined by the SHAP values. These different weight assignments can be attributed to GRU’s flexibility and capability to capture the business cycle asymmetries and nonlinearities. The SHAP method delivers some key recession indicators. For forecasting up to 3 months, the stock price index, real GDP, and private residential fixed investment show great short-term predictability, while for longer-term forecasting up to 12 months, the term spread and the producer price index have strong explanatory power for recessions. These findings are robust against other interpretation methods such as the local interpretable model-agnostic explanations (LIME) for GRU and the marginal effects for the ridge logit model.
@@ 1 Introduction
Recession forecasting has been a longstanding challenge for policymakers and market practitioners, as it enables them to make timely decisions that could mitigate the impact of a recession. However, due to the intricate and interconnected nature of modern economics, this task has proven to be difficult and has had limited success. Traditional linear methods such as probit or logit models, along with their variations and extensions, have been widely employed for this forecasting task. However, in recent decades, machine learning techniques, including artificial neural networks, have become increasingly popular among economists and have been applied to certain macroeconomic forecasting problems. Nevertheless, it has yet to be demonstrated whether and why these modern approaches allow for informed deviation from conventional linear models, when it comes to predicting recessions.
@@
Since the pioneering work of Mitchell and Burns (1938), which identified 21 variables out of a larger set as potential economic indicators for business cycles out of a larger set, researchers have been engaged in selecting indicators and developing theoretical frameworks or predictive models to link these indicators with business cycles. However, most studies have primarily focused on linear frameworks. Many of them have utilized probit regression models or their extensions to generate recession forecasts. Estrella and Mishkin (1996, 1998) compiled a combination of financial and macroeconomic variables and conducted recession forecasting using a probit framework. Their findings revealed that stock prices exhibit greater short-term predictability, while the slope of the yield curve performs better for longer-term predictions. Wright (2006) demonstrated that probit models incorporating the federal funds rate and term spread as predictors outperform models solely relying on the term spread. Dueker (1997, 2002) expanded the standard probit model by incorporating Markov regime switching within the probit framework, allowing for variation in coefficients. Chauvet and Potter (2005) introduced several specifications to the probit model that accounted for different business cycle dependencies and autocorrelation errors, concluding that their more complicated extensions improved the accuracy of recession forecasts. Fornari and Lemke (2010) as well as Nyberg (2014) incorporated vector autoregression components into the probit model to capture the endogenous dynamics of the predictors.
@@
While probit regression and its extensions continue to be widely used in business cycle forecasting, there has been a growing interest in exploring the forecasting capabilities of nonlinear models, including machine learning methods. This interest stems from the recognition that the business cycle often exhibits asymmetric and nonlinear patterns (Acemoglu and Scott (1997); Morley and Piger (2012)). Empirical evidence supports this notion, such as the work of Tiao and Tsay (1994), who demonstrate that a threshold autoregressive model outperforms a linear autoregressive model for predicting GDP growth. Maasoumi et al. (1994) examine multiple macroeconomic time series and confirm their nonlinear nature. Puglia and Tucker (2021) highlight the attractiveness of machine learning methods as an alternative to probit regression and its extensions, noting that probit methods typically require additional parameters for flexibility, whereas flexibility is inherent in machine learning methods. Stock and Watson (1998) compare the forecasting performance of 49 univariate linear and nonlinear models across 215 macroeconomic time series. They find that some of the nonlinear models perform poorly compared to linear models. Jaditz et al. (1998) explore the use of nearest neighbor regression models for forecasting industrial production but observe only marginal improvements in predictive performance. Vishwanathan and Murty (2002) present an iterative algorithm for support vector machines in classification problems. Ng (2014) applies a tree ensemble classifier to a large panel of predictors. Fornaro (2016) combines a Bayesian methodology with a shrinkage prior within the probit framework to predict recessions using extensive sets of predictors. More recently, Holopainen and Sarlin (2017), Bluwstein et al. (2020), and Vrontos et al. (2021) employ various machine learning methods for economic event forecasting. Vrontos et al. (2021) provide empirical evidence supporting the application of machine learning over traditional econometric techniques in the context of recession forecasting.
@@
Neural networks, as a subfield of machine learning, have garnered significant attention and application in fields like finance, primarily due to their ability to establish flexible mappings between variables and exhibit high pattern recognition capabilities (Zhang et al. (1998)). Neural networks, being highly nonlinear and nonparametric models, can approximate almost any functional form accurately, as stated by the universal approximation theorem (Hornik et al. (1989)), given that the network is wide or deep enough. Consequently, they are valuable modeling tools, particularly when there is limited prior knowledge about the appropriate functional relationships. However, the utilization of neural networks in macroeconomic studies has been relatively limited due to the small sample sizes and low-frequency nature of macroeconomic data. Swanson and White (1997) compare artificial neural networks with linear models in terms of predictive performance for nine macroeconomic variables, revealing only marginal improvements in forecast accuracy. Moshiri and Cameron (2000) apply neural networks to inflation forecasting using a dataset of 300 observations spanning 25 years of monthly data. Tkacz (2001) compares multivariate neural networks with linear and univariate models, finding minor forecast improvements in the short term but more pronounced benefits for longer horizons, such as a one-year forecast. Qi (2001) employs a simple feed-forward neural network to predict US recessions using a range of financial and economic indicators, identifying some indicators as useful for prediction. More recently, Puglia and Tucker (2021) compare neural networks with probit regression in forecasting US recessions using the term spread and other macro-financial variables, finding little difference between the models when evaluated. Wang et al. (2022) employ a specific type of recurrent neural network, namely a Bi-LSTM with autoencoder, along with other machine learning models to predict the beginning and end of economic recessions in the US. Their results suggest that the Bi-LSTM with autoencoder is the most accurate model.
@@
The novelty of this paper is twofold: Firstly, it focuses on two special types of recurrent neural networks, the long short-term memory (LSTM) and the gated recurrent unit (GRU), which address the limitations of a standard recurrent neural network related to the exploding and vanishing gradient problems. Their performance is compared to the simple feedforward neural network (FFN) that suffer from the key limitation of having to specify the temporal dependence upfront in the design of the model and to the traditional linear models in the context of recession forecasting. Secondly, the paper applies the Shapley additive explanations method (SHAP) to GRU which shows higher overall performance than LSTM, to explore the variable importance for different forecast horizons. In the context of recession forecasting, Puglia and Tucker (2021) and Delgado et al. (2022) also use the SHAP method to decompose recession forecasts, but they are applied to other models than LSTM and GRU. The three main findings can be summarized as follows: Firstly, the out-of-sample performance strongly supports the application of LSTM and GRU in the area of recession forecasting, especially for the long-term forecasting tasks. They outperform other types of models across five forecast horizons with respect to different types of statistical performance metrics. Secondly, GRU and the ridge logit model differ in assessing variable importance, evident in the different variable orders based on the SHAP values. Lastly, while the leading predictors for GRU and ridge logit models slightly differ, key indicators like S&P 500 index, real GDP, and private residential fixed investment consistently emerge for short-term predictions (up to 3 months). For longer-term forecasts (6 months or more), the term spread and producer price index take precedence. These results are corroborated by local interpretable model-agnostic explanations (LIME) and marginal effects, respectively.
@@ 2 Data
Prior studies in business cycle forecasting often rely on macroeconomic indicators, which are subject to revisions after their initial estimates. Stark and Croushore (2002) demonstrate that the accuracy of forecasts is influenced by using the most up-to-date data instead of real-time data. Therefore, when comparing forecasts from new models to benchmark forecasts, it is crucial to ensure that the comparisons are based on real-time data. In my research, the focus is primarily on assessing the real-time predictability of neural network models with respect to the Great Recession and the Covid-19 recession in the United States. This necessitates working with real-time data since the information that is available in hindsight was not accessible prior to the recessions. To evaluate the predictability of the recessions, I utilize the same data that actually was available to real-time forecasters for out-of-sample forecasting.
The dataset employed for prediction consists of 194 real-time vintages of macroeconomic and financial market variables, covering the period from February 1967 to October 2021. The out-of-sample forecasting commences in November 2006, utilizing the real-time vintage of data. Taking into account previous studies like Vrontos et al. (2021) and the availability of real-time data, a set of 25 predictors is chosen. A detailed list with descriptions of these variables can be found in Table 1.
@@ 25 Real personal income excluding current transfer receipts W875RX1 Income Log growth rate Monthly
The selected predictors cover a wide range of categories, including output, income, prices, the labor market, the housing market, money and credit, and the financial market. The data frequency varies from daily to quarterly. To facilitate model estimation, higher frequency data is aggregated into monthly data using the mean. Quarterly frequency variables are transformed into monthly equivalents using natural cubic spline interpolation. Specifically, at each month, all available data up to that point are used to calculate the interpolating cubic spline. This spline curve is then utilized to generate data of monthly frequency that lie on the curve between the quarterly data points.
The majority of monthly data vintages for the variables are obtained from Archival Federal Reserve Economic Data (ALFRED). However, there are a few selected predictors for which the real-time data prior to 2013 is not available in ALFRED. The first set of variables, including real personal income excluding transfer receipts and real manufacturing and trade sales, along with total non-farm payroll employment and the industrial production index, are used by Chauvet and Piger (2008) to identify business cycle dates in real time. The real-time data for these series, provided by Jeremy Piger on his website, ends in August 2013. However, it can be easily extended beyond 2013 by using the data in ALFRED. The second set of variables consists of real M1 and M2 money stock, for which the earliest available real-time data in ALFRED is from January 2014. To address the absence of real-time vintages for real M1 and M2 money stock before this date, nominal M1 and M2 money stock are adjusted for inflation using the consumer price index, which has real-time vintages readily accessible in ALFRED.
To identify recession periods in the United States, I rely on the business cycle expansion and contraction dates determined by the National Bureau of Economic Research (NBER). NBER is widely regarded as the standard reference for US business cycles in the existing literature. In this study, recession months are defined as the period following the peak and continuing until the trough, while all other months are considered periods of economic expansion. The earliest available vintage of NBER recession indicator in ALFRED is from September 2014. For the monthly vintages preceding that date, I manually collect and construct them based on the official announcements made by the NBER business cycle dating committee [^1] https://www.nber.org/research/business-cycle-dating/business-cycle-dating-committee-announcements. However, it’s worth noting that one major practical challenge with NBER business cycle dates is that they are often announced with significant publication delays.
@@
The table reports NBER business cycle dates, including the type of cycle (contraction or expansion), the duration in months, and the time of announcement. The data covers the period from 1980 to 2021, and publication lags in months are indicated in parentheses.
Table 2 presents the peak and trough dates of the US business cycle from 1980 to 2021, along with their corresponding announcement dates. The publication lags for the recessions listed in the table range from 5 to 21 months, with troughs being identified later than peaks on average. While the NBER business cycle dates remain unchanged once finalized, the presence of publication lags complicates the creation of real-time versions of the NBER recession indicator. To address this, the NBER recession indicator in ALFRED is constructed under the assumption that the previous state remains unchanged until a new turning point is officially announced.
@@ 3 Econometric methodology
In this section, the focus is on the technical aspects of the models used to predict the two most recent recessions in the United States. The paper specifically explores neural networks, which are complex nonlinear models composed of interconnected nodes arranged in multiple layers. These networks have the ability to approximate any linear or nonlinear continuous functions, as stated by the universal approximation theorem (Hornik et al. (1989)), given that the network is wide or deep enough. In this paper, three types of neural networks are employed for recession forecasting: feedforward neural network (FFN), long short-term memory (LSTM), and gated recurrent unit (GRU). Section 3.1 provides a detailed description of these models. Model specifications, including estimation and prediction techniques, are discussed in Section 3.2. Finally, Section 3.3 introduces statistical measures that effectively evaluate the prediction performance.
@@ 3.1.1 Feedforward neural network
FFN is a widely used and straightforward type of artificial neural network. It consists of multiple processing units called nodes or neurons that are organized into layers. It operates by transmitting information in a unidirectional manner, where data flows from the input layer to the output layer without feedback loops. Figure 1 illustrates, based on the NN-SVG  [^2] Lenail (2019) https://doi.org/10.21105/joss.00747, an example of a three-layer FFN designed for binary classification. It comprises an input layer with eight units, a hidden layer with four units, and an output layer with one unit. The first units in the input and hidden layers serve as bias units. In this configuration, data from the input layer is passed through the hidden layer, which transforms the data. The values obtained from the hidden layer are then forwarded to the output layer, which translates them into desired outputs based on the problem at hand. In the case of binary classification, the last unit in the output layer utilizes a sigmoid activation function, producing a value between 0 and 1. This value represents the probability of an event occurring.
Depending on the number of explanatory variables in the data and the number of units in the hidden layer the unknown underlying function $f$ for an output node can be written as
@@
where $n$ is the number of predictors, $k$ is the number of units in the hidden layer, $g_{1}$ is the activation function in the hidden layer, $g_{2}$ is the activation function in the output layer, $\beta_{ij}$ and $\alpha_{j}$ represent the weight parameters from the input to the hidden layer and from the hidden to the output layer, respectively, and $\epsilon$ is the error term. The weight parameters are estimated by the backpropagation process that seeks to repeatedly update the weights until convergence based on the derivatives of the cost function with respect to input and the hidden layers.

### v3

@@ Abstract
Long short-term memory (LSTM) and gated recurrent unit (GRU) are used to model US recessions from 1967 to 2021. Their predictive performances are compared to those of the traditional linear models. The out-of-sample performance suggests the application of LSTM and GRU in recession forecasting, especially for longer-term forecasts. The Shapley additive explanations (SHAP) method is applied to both groups of models. The SHAP-based different weight assignments imply the capability of these types of neural networks to capture the business cycle asymmetries and nonlinearities. The SHAP method delivers key recession indicators, such as the S&P 500 index for short-term forecasting up to 3 months and the term spread for longer-term forecasting up to 12 months. These findings are robust against other interpretation methods, such as the local interpretable model-agnostic explanations (LIME) and the marginal effects.
@@ 1 Introduction
Recession forecasting has been a longstanding challenge for policymakers and market practitioners, enabling them to make timely decisions that could mitigate the impact of a recession. However, due to the interconnected nature of modern economics, this task has proven difficult and has had limited success. Treating recession as a binary event, traditional linear methods such as probit and logit models and extensions have been widely employed for this forecasting task. However, in recent decades, machine learning techniques, including artificial neural networks, have become increasingly popular among economists and have been applied to certain macroeconomic forecasting problems. Nevertheless, whether and why these modern approaches allow for informed deviation from conventional linear models when predicting recessions has yet to be demonstrated.
@@
Since the pioneering work of Mitchell and Burns (1938), which identified 21 variables out of a larger set as potential economic indicators for business cycles, researchers have been engaged in selecting indicators and developing theoretical frameworks and predictive models to link these indicators with business cycles. However, most studies have primarily focused on linear frameworks. Many have used probit regression models and extensions to generate recession forecasts. Estrella and Mishkin (1996, 1998) compiled a combination of financial and macroeconomic variables and conducted recession forecasting using a probit framework. Their findings revealed that stock prices exhibit greater short-term predictability, while the yield curve slope performs better for longer-term predictions. Wright (2006) demonstrated that probit models incorporating the federal funds rate and term spread as predictors outperform models solely relying on the term spread. Dueker (1997, 2002) expanded the standard probit model by incorporating Markov regime switching within the probit framework, allowing for coefficient variation. Chauvet and Potter (2005) introduced several specifications to the probit model that accounted for different business cycle dependencies and autocorrelation errors, concluding that their more complicated extensions improved the accuracy of recession forecasts. Fornari and Lemke (2010) and Nyberg (2014) incorporated vector autoregression components into the probit model to capture the endogenous dynamics of the predictors.
@@
While probit regression and its extensions are widely used in business cycle forecasting, there has been a growing interest in exploring the predictability of nonlinear models, including machine learning methods. This interest stems from recognizing that the business cycle often exhibits asymmetric and nonlinear patterns (Acemoglu and Scott (1997); Morley and Piger (2012)). Empirical evidence supports this notion, such as the work of Tiao and Tsay (1994), who demonstrate that a threshold autoregressive model outperforms a linear autoregressive model for predicting GDP growth. Maasoumi et al. (1994) examine multiple macroeconomic time series and confirm their nonlinear nature. Puglia and Tucker (2021) highlight the attractiveness of machine learning methods as an alternative to probit regression and its extensions, noting that probit methods typically require additional parameters for flexibility. In contrast, flexibility is inherent in machine learning methods. Stock and Watson (1998) compare the forecasting performance of 49 univariate linear and nonlinear models across 215 macroeconomic time series. They find that some nonlinear models perform poorly compared to linear models. Jaditz et al. (1998) explore using nearest neighbor regression models for forecasting industrial production but observe only marginal improvements in predictive performance. Vishwanathan and Murty (2002) present an iterative algorithm for support vector machines in classification problems. Ng (2014) applies a tree ensemble classifier to a large panel of predictors. Fornaro (2016) combines a Bayesian methodology with a shrinkage prior within the probit framework to predict recessions using extensive predictors. More recently, Holopainen and Sarlin (2017), Bluwstein et al. (2020), and Vrontos et al. (2021) employ various machine learning methods for economic event forecasting. Vrontos et al. (2021) provide empirical evidence supporting the application of machine learning over traditional econometric techniques in recession forecasting.
@@
As a subfield of machine learning, neural networks have gained significant attention and application in fields like finance, primarily due to high pattern recognition capabilities and their ability to establish flexible mappings between the variables (Zhang et al. (1998)). Neural networks, being highly nonlinear and nonparametric, can approximate almost any functional form accurately, as stated by the universal approximation theorem (Hornik et al. (1989)), given that the network is wide or deep enough. Consequently, they are valuable modeling tools, particularly when there is limited prior knowledge about the appropriate functional relationships. However, using neural networks in macroeconomic studies has been relatively limited due to the small sample sizes and low-frequency nature of macroeconomic data. Swanson and White (1997) compare artificial neural networks with linear models regarding predictive performance for nine macroeconomic variables, revealing only marginal improvements in forecast accuracy. Moshiri and Cameron (2000) apply neural networks to inflation forecasting using a dataset of 300 observations spanning 25 years of monthly data. Tkacz (2001) compares multivariate neural networks with linear and univariate models, finding minor forecast improvements in the short term but more pronounced benefits for longer horizons, such as a one-year forecast. Qi (2001) employs a simple feed-forward neural network to predict US recessions using a range of financial and economic indicators, identifying some indicators as useful for prediction. More recently, Puglia and Tucker (2021) compare neural networks with probit regression in forecasting US recessions using the term spread and other macro-financial variables, finding little difference between the models when evaluated. Wang et al. (2022) employ a specific type of recurrent neural network, namely a Bi-LSTM with autoencoder, along with other machine learning models to predict the beginning and end of economic recessions in the US. Their results suggest that the Bi-LSTM with autoencoder is the most accurate model.
@@
The novelty of this paper is twofold: Firstly, it focuses on two special types of recurrent neural networks, the long short-term memory (LSTM) and the gated recurrent unit (GRU), which address the limitations of a standard recurrent neural network related to the exploding and vanishing gradient problems. Their performance is compared to the simple feedforward neural network (FFN), which suffers from the key limitation of specifying the temporal dependence upfront in the model’s design and to the traditional linear models in the context of recession forecasting. Secondly, the paper applies the Shapley additive explanations method (SHAP) to GRU, which shows higher overall performance than LSTM, to explore the variable importance of different forecast horizons. In recession forecasting, Puglia and Tucker (2021) and Delgado et al. (2022) also use the SHAP method to decompose recession forecasts, but they are applied to models other than LSTM and GRU. The three main findings can be summarized as follows: Firstly, the out-of-sample performance strongly supports the application of LSTM and GRU in recession forecasting, especially for long-term forecasting tasks. They outperform other types of models across five forecast horizons for different types of statistical performance metrics. Secondly, GRU and the ridge logit model differ in assessing variable importance, evident in the different variable orders based on the SHAP values. Lastly, while the leading predictors for GRU and ridge logit models slightly differ, key indicators like S&P 500 index, real GDP, and private residential fixed investment consistently emerge for short-term predictions (up to 3 months). The term spread and producer price index precede in longer-term forecasts (6 months or more). These results are corroborated by local interpretable model-agnostic explanations (LIME) and marginal effects.
@@ 2 Data
Prior studies in business cycle forecasting often rely on macroeconomic indicators, subject to revisions after initial estimates. Stark and Croushore (2002) demonstrate that the accuracy of forecasts is influenced by using the most up-to-date data instead of real-time data. Therefore, when comparing forecasts from new models to benchmark forecasts, it is crucial to ensure that the comparisons are based on real-time data. My research focuses primarily on assessing the real-time predictability of neural network models for the Great Recession and the COVID-19 recession in the United States. This necessitates working with real-time data since the information available in hindsight was inaccessible before the recession. To evaluate the predictability of the recessions, I utilized the same data available to real-time forecasters for out-of-sample forecasting.
The dataset employed for prediction consists of 194 real-time vintages of macroeconomic and financial market variables, covering the period from February 1967 to October 2021. The out-of-sample period begins in November 2006. Considering previous studies and real-time data availability, a set of 25 predictors is chosen. A detailed list with descriptions of these variables can be found in Table 1.
@@ 25 Real personal income excluding current transfer receipts W875RX1 Income Log growth rate Monthly
The selected predictors cover various categories: output, income, prices, labor, housing, money and credit, and financial markets. The data frequency varies from daily to quarterly. Higher frequency data is aggregated into monthly data using the mean. Quarterly frequency variables are transformed into monthly equivalents using natural cubic spline interpolation. Specifically, at each month, all available data up to that point are used to calculate the interpolating cubic spline. This spline curve is then utilized to generate data of monthly frequency that lie on the curve between the quarterly data points.
Most monthly data vintages for the variables are obtained from Archival Federal Reserve Economic Data (ALFRED). However, there are two groups of variables for which the real-time data before 2013 is not available in ALFRED. The first set of variables, including real personal income excluding transfer receipts, real manufacturing and trade sales, total non-farm payroll employment, and the industrial production index, are used by Chauvet and Piger (2008) to identify business cycle dates in real time. The real-time data for these series, provided by Jeremy Piger on his website, ends in August 2013. However, it can be easily extended beyond 2013 using the ALFRED data. The second set of variables consists of real M1 and M2 money stock, for which the earliest available real-time data in ALFRED is from January 2014. To address the absence of real-time vintages for real M1 and M2 money stocks before this date, nominal M1 and M2 money stocks are adjusted for inflation using the consumer price index, which has real-time vintages readily accessible in ALFRED.
To identify recession periods in the United States, I rely on the business cycle expansion and contraction dates determined by the National Bureau of Economic Research (NBER). NBER is the standard reference for US business cycles in the existing literature. This study defines recession months as the period following the peak and continuing until the trough, while all other months are considered periods of economic expansion. The earliest available vintage of the NBER recession indicator in ALFRED is September 2014. For the monthly vintages preceding that date, I manually collect and construct them based on the official announcements made by the NBER business cycle dating committee [^1] https://www.nber.org/research/business-cycle-dating/business-cycle-dating-committee-announcements.
@@
The table reports NBER business cycle dates, including the type of cycle (contraction or expansion), the duration in months, and the time of announcement. The data covers the period from 1980 to 2021; publication lags in months are indicated in parentheses.
One major practical challenge with NBER business cycle dates is that they are often announced with significant publication delays. Table 2 presents the peak and trough dates of the US business cycle from 1980 to 2021, along with their corresponding announcement dates. The publication lags for the recessions in the table range from 5 to 21 months, with troughs being identified later than peaks on average. While the NBER business cycle dates remain unchanged once finalized, publication lags complicate the creation of real-time versions of the NBER recession indicator. To address this, the NBER recession indicator in ALFRED is constructed assuming the previous state remains unchanged until a new turning point is officially announced.
@@ 3 Econometric methodology
This section focuses on the technical aspects of the models used to predict the two most recent recessions in the United States. The paper explores neural networks, complex nonlinear models of interconnected nodes arranged in multiple layers. Given that the network is wide or deep enough, these networks can approximate any linear or nonlinear continuous functions, as stated by the universal approximation theorem (Hornik et al. (1989)). In this paper, three types of neural networks are employed for recession forecasting: feedforward neural network (FFN), long short-term memory (LSTM), and gated recurrent unit (GRU). Section 3.1 provides a detailed description of these models. Model specifications, including estimation and prediction techniques, are discussed in Section 3.2. Finally, Section 3.3 introduces statistical measures that effectively evaluate the prediction performance.
@@ 3.1.1 Feedforward neural network
FFN is a widely used and straightforward type of artificial neural network. It consists of multiple processing units called nodes or neurons organized into layers. It operates by transmitting information in a unidirectional manner, where data flows from the input layer to the output layer without feedback loops. Figure 1 illustrates, based on the NN-SVG  [^2] Lenail (2019) https://doi.org/10.21105/joss.00747, an example of a three-layer FFN designed for binary classification. It comprises an input layer with eight units, a hidden layer with four units, and an output layer with one unit. The first units in the input and hidden layers serve as bias units. In this configuration, data from the input layer is passed through the hidden layer, which transforms the data. The values obtained from the hidden layer are then forwarded to the output layer, translating into desired outputs based on the problem. In the case of binary classification, the last unit in the output layer utilizes a sigmoid activation function, producing a value between 0 and 1. This value represents the probability of an event occurring.
Given this architecture, the unknown underlying function $f$ for an output node can be written as
@@
where $n$ is the number of predictors, $k$ is the number of units in the hidden layer, $g_{1}$ is the activation function in the hidden layer, $g_{2}$ is the activation function in the output layer, $\beta_{ij}$ and $\alpha_{j}$ represent the weight parameters from the input to the hidden layer and from the hidden to the output layer, respectively, and $\epsilon$ is the error term. The backpropagation process estimates the weight parameters, which seeks to repeatedly update the weights until convergence based on the derivatives of the cost function for input and the hidden layers.